# NeoVerse: Enhancing 4D World Model with in-the-wild Monocular Videos

### v1

@@ 1 Introduction
4D world modeling holds transformative potential in many fields, such as digital content creation, autonomous driving, and embodied intelligence. Recent approaches have made strides from both 3D side [gen3c, difix3d, viewcrafter, voyager, see3d, flexworld] and 4D side [trajectorycrafter, ex4d, uni3c, free4d, freesim, flexdrive, vivid4d, chronosobserver, ma2025follow, worldforge, postcam, van2024generative, cinemaster] with a principle of hybrid reconstruction and generation. This paradigm typically involves two stages: reconstructing a 3D/4D representation [depthcrafter, yang2024depth, bochkovskii2024depth, 3dgs, monst3r] of the scene, and then, using the geometric prior to guide generation models [stable_diffusion, cogvideo, cogvideox, hunyuan_dit, wan]. Such a reconstruction-generation hybrid paradigm has widely recognized promising features, including spatiotemporal consistency and precise viewpoint control. However, the current solutions usually have limitations in terms of scalability.
@@
The limitation of scalability manifests in two main aspects. (1) Limited data scalability. Some methods, such as ViewCrafter [viewcrafter], utilize videos of static scenes to create multi-view training data and learn to generate videos in novel trajectories. Although effective, they cannot be extended to 4D scenes. Some other methods, such as SynCamMaster [syncammaster], CamCloneMaster [camclonemaster], and ReCamMaster [recammaster], depend on specialized, hard-to-capture multi-view dynamic videos to learn novel trajectory generation. Such non-scalable data limits the model’s generalization and versatility. (2) Limited training scalability. Another line of work [trajectorycrafter, freesim, uni3c, vivid4d] utilizes more flexible data types but usually necessitates a cumbersome offline pre-processing stage to create training data. For example, TrajectoryCrafter generates training data using a heavy video depth estimator [depthcrafter] in an offline manner. Similarly, previous work FreeSim [freesim] pre-reconstructs the Gaussian field to prepare training input, which utilizes offline reconstruction [pvg, omnire] and may even rely on extra 3D detection methods [centerpoint, bevformer, sst, mixsup]. Such an offline curation usually leads to significant computational burden, storage consumption, inflexible training scheme tuning, and even disables online augmentations. The two kinds of limitations erect a barrier to leveraging the cheap and diverse in-the-wild monocular videos, constraining the potential for building more powerful models.
@@
To address these challenges, we propose NeoVerse. The core philosophy of NeoVerse is making the full pipeline scalable to diverse in-the-wild monocular videos, enhancing generalization and versatility of 4D world models. To implement our vision, we first propose a feed-forward 4DGS model, built upon VGGT [vggt]. This model not only “Gaussianizes” VGGT but also features a bidirectional motion modeling mechanism, which is crucial for efficient online reconstruction (Sec. 3.2) and applications requiring time control. We then incorporate this feed-forward model into the generation training process. During each training iteration, it efficiently reconstructs 4D scenes using sparse key frames from monocular videos in an online manner. In addition, efficient online monocular degradation simulations, including Gaussian culling and average geometry filter, are proposed to simulate degraded rendering patterns in novel trajectories and offer conditions for generation. Combining them together makes the whole training process scalable to diverse in-the-wild monocular videos (up to 1M clips) in terms of both training efficiency and technical feasibility. We summarize our contributions as follows.
@@ 2 Related Works
Recent stereo and 3D geometry foundation models [dust3r, noposplat, monst3r, vggt, anysplat, mapanything, worldmirror, da3, 4dgt, streamsplat, movies] can estimate dense depth, point maps, and even camera parameters in a single forward pass, thereby driving a shift in Gaussian Splatting from per-scene optimization to generalizable feed-forward reconstruction. For static scenes, pose-free models such as NoPoSplat [noposplat] reconstruct 3D Gaussians directly from sparse, unposed multi-view images, and AnySplat [anysplat] further extends this paradigm to casually captured, long uncalibrated image sequences. For dynamic scenes, 4DGT [4dgt], StreamSplat [streamsplat], and MoVieS [movies] push feed-forward GS into 4D; however, each method still retains specific constraints: 4DGT is trained on posed monocular videos and adopts a largely uni-directional temporal modeling strategy, MoVieS similarly assumes known camera poses during training and inference, while StreamSplat focuses on frame-by-frame modeling.
@@
Recent methods such as GEN3C [gen3c], DaS [das], See3D [see3d], ViewCrafter [viewcrafter], Difix3D+ [difix3d], GS-DiT [gsdit], Voyager [voyager], Uni3C [uni3c], FreeSim [freesim], TrajectoryCrafter [trajectorycrafter], See4D [see4d], PostCam [postcam], Light-X [lightx] follow a hybrid reconstruction+generation paradigm, where a 3D/4D representation is first reconstructed and then used as geometric guidance for a generative video model. GEN3C [gen3c] builds a depth-based 3D feature cache whose renderings condition a video diffusion model for 3D-consistent, pose-controllable synthesis; ViewCrafter [viewcrafter] adopts a point-conditioned video diffusion framework to extend single- or sparse-view inputs into long-range, high-fidelity novel-view sequences; Difix3D+ [difix3d] applies a single-step diffusion enhancer to rendered novel views to correct artifacts in underconstrained regions and distill the improvements back into NeRF/3DGS representations; and TrajectoryCrafter [trajectorycrafter] formulates camera-controllable video generation for monocular videos as trajectory redirection, conditioning a dual-stream diffusion backbone on point-cloud renderings and source frames to follow user-specified camera paths. Despite their strong spatial–temporal consistency and viewpoint controllability, these reconstruction-based approaches are mostly tailored to static or quasi-static scenes and rely on curated data or heavyweight offline reconstruction, limiting scalability to in-the-wild monocular videos.
@@ 3.1 Pose-Free Feed-Forward 4DGS Reconstruction
Our feed-forward model is partially built upon VGGT [vggt] backbone. For simplicity, we mainly introduce how we make VGGT dynamic and “Gaussianized”.
@@
Given a monocular video $\{\boldsymbol{I}_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=1}^{T}$, VGGT extracts the frame-wise features using the pretrained DINOv2 [dinov2]. These features, concatenated with camera tokens and register tokens, are fed into a series of Alternating-Attention blocks [vggt], obtaining so-called frame features. While this process effectively aggregates spatial information, they are insufficient for motion modeling due to temporal unawareness.
We introduce a bidirectional motion-encoding branch. Different from uni-directional motion in 4DGT [4dgt], the bidirectional prediction distinguishes the instantaneous velocity between $t\rightarrow t+1$ and $t\rightarrow t-1$. Such a distinction facilitates temporal Gaussian interpolation between two consecutive timestamps.
@@
where each Gaussian $i$ is parameterized by: 3D position $\boldsymbol{\mu}_{i}$, opacity $\boldsymbol{\alpha}_{i}$, rotation $\boldsymbol{r}_{i}$, scale $\boldsymbol{s}_{i}$, and spherical harmonics coefficients $\boldsymbol{sh}_{i}$, as inherited from 3D Gaussians [3dgs]. For bidirectional motion modeling, we introduce forward and backward velocities $\boldsymbol{v}^{+}_{i}$, $\boldsymbol{v}^{-}_{i}$, and forward and backward angular velocities $\boldsymbol{\omega}^{+}_{i}$, $\boldsymbol{\omega}^{-}_{i}$. In addition, we adopt a life span $\boldsymbol{\tau}_{i}$ following the common practice in 4DGS.
@@ 3.2 Reconstruction-guided Video Generation
Our generation model is expected to generate high-quality novel views from low-quality novel view renderings, necessitating such training pairs. For multi-view or static datasets [scannet++, dl3dv], we can easily get such training pairs as in ViewCrafter [viewcrafter]. However, for in-the-wild monocular videos, we need to carefully simulate degradation renderings paired with ground-truth monocular frames. Therefore, we propose three techniques to simulate the degradation rendering patterns based on monocular videos.
@@
(2) Average Geometry Filter for flying-edge-pixel and distortion simulation. In addition to the occlusion, another typical degradation pattern is the flying pixels in depth-discontinuous edges. The network has tendency to produce average depth value at those edges to minimize regression loss, as also confirmed by [pixelperfect]. From a first-principles perspective, we propose to use a average filter to create such averaged depth patterns. Specifically, we render depth in the transformed novel trajectory and apply an average filter in the rendered depth map. We then adjust the center position of each Gaussian according to the average filtered depth value. When such modified Gaussians are rendered back into the original views, the flying-pixel pattern appears as shown in Fig. 3 (b). We can further apply a larger filter kernel to simulate spatially broader distortions shown in Fig. 3 (c), caused by potential depth error.
@@
We use the obtained degraded renderings as conditions for generation and the original videos as targets. The rendered conditions include multiple modalities, including RGB images, depth maps, and masks binarized from opacity maps to indicate the empty regions. Plüker embeddings of the original trajectory are also computed to provide explicit 3D camera motion information [uni3c]. We introduce a control branch to incorporate them into the generation model like [controlnet, vace, voyager, layeranimate]. During training, we only train the control branch while freezing the video generation model, not only for training efficiency, but more importantly, to make NeoVerse accessible to powerful distillation LoRAs [lora] to speed up the generation process.
@@ 3.3 Training Scheme
where $\mathcal{L}_{\text{rgb}}$ is the photometric loss between rendered and ground-truth images, including an $L_{2}$ loss and LPIPS [lpips] loss. The camera loss $\mathcal{L}_{\text{camera}}$ and depth loss $\mathcal{L}_{\text{depth}}$ supervise the predicted camera parameters and depth maps following VGGT [vggt]. Notably, $\mathcal{L}_{\text{depth}}$ also contains the supervision for rendered depth from Gaussians. The motion loss $\mathcal{L}_{\text{motion}}=\sum_{i}\|\hat{\boldsymbol{v}}^{+}_{i}-\boldsymbol{v}^{+}_{i}\|+\|\hat{\boldsymbol{v}}^{-}_{i}-\boldsymbol{v}^{-}_{i}\|$ adds supervision on the predicted bidirectional velocities, where $\hat{\boldsymbol{v}}^{+}_{i}$ and $\hat{\boldsymbol{v}}^{-}_{i}$ are the ground-truth forward and backward velocities computed from some dynamic 3D datasets [pointodyssey, dynamicstereo, kubric, spring, waymo, vkitti2]. To prevent the Gaussians from becoming erroneously transparent, we introduce a regularization loss $\mathcal{L}_{\text{regular}}=\sum_{i}|1-\boldsymbol{A}_{i}|$, where $\boldsymbol{A}_{i}$ is rendered accumulated opacity map.
@@
For generation model training, we adopt Rectified Flow [rectified] and Wan-T2V [wan] 14B to model the denoising diffusion process. The whole training process is performed on monocular videos. Given a monocular video, we first utilize on-the-fly reconstruction from sparse key frames to obtain 4DGS and simulate degradation renderings as conditions $c_{\text{render}}$. For the video latent $x_{1}$ and sampled noise $x_{0}\sim\mathcal{N}(0,I)$, the training objective of generation model $f_{\theta}$ is formulated as
@@
where $x_{t}$ is a linear interpolation between $x_{1}$ and $x_{0}$ at timestamp $t$, $v_{t}=x_{1}-x_{0}$ is ground-truth velocity. $c_{\text{text}}$ is the text condition extracted from the video caption using a language model like umT5 [umT5]. Renderings $c_{\text{render}}$ are input into the generation model through a control branch like [controlnet, vace].
@@ 4.1 Implementation
For reconstruction, we follow the learning rate schedule of VGGT [vggt]. We resize all input videos to have a longest edge of 560 pixels. GSplat [gsplat] is adopted as the Gaussian Splatting rendering backend. For the generation, the video resolution is fixed at $336\times 560$ and the length is set to $81$ frames. The training is conducted on 32 A800 GPUs, where the first stage trains 150K iterations and the second stage trains 50K iterations. More training details can be found in the supplementary material.
@@
We collect 18 public datasets following CUT3R [cut3r], including Arkitscenes [arkitscenes], DL3DV [dl3dv], PointOdyssey [pointodyssey], Kubric [kubric], Waymo [waymo], SpatialVID [spatialvid], GFIE [gfie], etc. Besides the above datasets, we further curate a large-scale self-collected monocular video dataset from the internet, containing over 1M videos from diverse scenarios. More details about datasets are provided in the supplementary material.
@@ 4.2 Quantitative Evaluation
Our reconstruction results on both static and dynamic datasets are shown in Table 1 and Table 2, respectively. Our reconstruction part achieves state-of-the-art performance among all metrics. Recent reprints MoVieS [movies] and StreamSplat [streamsplat] are not listed in the table because they are neither open-sourced nor provide a detailed evaluation protocol. Our detailed evaluation protocols are provided in the supplementary material.
Method VRNeRF [vrnerf] (16 views) Scannet++ [scannet++] (32 views)
@@
NoPoSplat [noposplat] 11.27 0.408 0.620 8.69 0.312 0.614
Flare [flare] 12.62 0.597 0.623 12.19 0.619 0.611
AnySplat [anysplat] 18.02 0.705 0.366 22.79 0.773 0.217
@@
In Table 3, we compare the generation performance with related work TrajectoryCrafter [trajectorycrafter] and ReCamMaster [recammaster], demonstrating better performance. We conduct more analysis in the section of qualitative evaluation.
@@
Table 3 also shows the efficiency evaluation of both the reconstruction stage and the generation stage. Thanks to our intentional design of condition injection in Sec. 3.2, our generation process gets significantly accelerated by the off-the-shelf distillation technique [lightx2v]. More importantly, as discussed in Sec. 3.2, our bidirectional motion design enables more efficient reconstruction from sparse key frames without loss of generation performance.
@@
TrajectoryCrafter [trajectorycrafter] 49 38 121 159 83.02 88.58 94.71 97.64 44.63 54.59
ReCamMaster [recammaster] 81 - 168 168 88.21 91.60 96.56 98.86 44.29 58.87
@@
Table 3: VBench [vbench] results for novel view generation. We randomly collect 100 unseen in-the-wild videos, each with 4 different camera trajectories, resulting in a total of 400 test cases. For a fair comparison of inference time, we resize all videos to $336\times 560$ resolution and report the average results over all test cases. The runtime evaluation is conducted on an A800 GPU.

### v2

@@ 1 Introduction
4D world modeling holds transformative potential in many fields, such as digital content creation, autonomous driving, and embodied intelligence. Recent approaches have made strides from both 3D side [58, 73, 88, 30, 51, 10] and 4D side [87, 25, 9, 46, 18, 94, 28, 70, 52, 62, 11, 64, 69] with a principle of hybrid reconstruction and generation. This paradigm typically involves two stages: reconstructing a 3D/4D representation [26, 79, 7, 36, 89] of the scene, and then, using the geometric prior to guide generation models [60, 23, 82, 40, 65]. Such a reconstruction-generation hybrid paradigm has widely recognized promising features, including spatiotemporal consistency and precise viewpoint control. However, the current solutions usually have limitations in terms of scalability.
@@
The limitation of scalability manifests in two main aspects. (1) Limited data scalability. Some methods, such as ViewCrafter [88], utilize videos of static scenes to create multi-view training data and learn to generate videos in novel trajectories. Although effective, they cannot be extended to 4D scenes. Some other methods, such as SynCamMaster [3], CamCloneMaster [50], and ReCamMaster [2], depend on specialized, hard-to-capture multi-view dynamic videos to learn novel trajectory generation. Such non-scalable data limits the model’s generalization and versatility. (2) Limited training scalability. Another line of work [87, 18, 9, 28] utilizes more flexible data types but usually necessitates a cumbersome offline pre-processing stage to create training data. For example, TrajectoryCrafter generates training data using a heavy video depth estimator [26] in an offline manner. Similarly, previous work FreeSim [18] pre-reconstructs the Gaussian field to prepare training input, which utilizes offline reconstruction [12, 13] and may even rely on extra 3D detection methods [86, 41, 17, 81]. Such an offline curation usually leads to significant computational burden, storage consumption, inflexible training scheme tuning, and even disables online augmentations. The two kinds of limitations erect a barrier to leveraging the cheap and diverse in-the-wild monocular videos, constraining the potential for building more powerful models.
@@
To address these challenges, we propose NeoVerse. The core philosophy of NeoVerse is making the full pipeline scalable to diverse in-the-wild monocular videos, enhancing generalization and versatility of 4D world models. To implement our vision, we first propose a feed-forward 4DGS model, built upon VGGT [67]. This model not only “Gaussianizes” VGGT but also features a bidirectional motion modeling mechanism, which is crucial for efficient online reconstruction (Sec. 3.2) and applications requiring time control. We then incorporate this feed-forward model into the generation training process. During each training iteration, it efficiently reconstructs 4D scenes using sparse key frames from monocular videos in an online manner. In addition, efficient online monocular degradation simulations, including Gaussian culling and average geometry filter, are proposed to simulate degraded rendering patterns in novel trajectories and offer conditions for generation. Combining them together makes the whole training process scalable to diverse in-the-wild monocular videos (up to 1M clips) in terms of both training efficiency and technical feasibility. We summarize our contributions as follows.
@@ 2 Related Works
Recent stereo and 3D geometry foundation models [71, 83, 89, 67, 32, 35, 47, 43, 78, 74, 42] can estimate dense depth, point maps, and even camera parameters in a single forward pass, thereby driving a shift in Gaussian Splatting from per-scene optimization to generalizable feed-forward reconstruction. For static scenes, pose-free models such as NoPoSplat [83] reconstruct 3D Gaussians directly from sparse, unposed multi-view images, and AnySplat [32] further extends this paradigm to casually captured, long uncalibrated image sequences. For dynamic scenes, 4DGT [78], StreamSplat [74], and MoVieS [42] push feed-forward GS into 4D; however, each method still retains specific constraints: 4DGT is trained on posed monocular videos and adopts a largely uni-directional temporal modeling strategy, MoVieS similarly assumes known camera poses during training and inference, while StreamSplat focuses on frame-by-frame modeling.
@@
Recent methods such as GEN3C [58], DaS [22], See3D [51], ViewCrafter [88], Difix3D+ [73], GS-DiT [5], Voyager [30], Uni3C [9], FreeSim [18], TrajectoryCrafter [87], See4D [49], PostCam [11], Light-X [45] follow a hybrid reconstruction+generation paradigm, where a 3D/4D representation is first reconstructed and then used as geometric guidance for a generative video model. GEN3C [58] builds a depth-based 3D feature cache whose renderings condition a video diffusion model for 3D-consistent, pose-controllable synthesis; ViewCrafter [88] adopts a point-conditioned video diffusion framework to extend single- or sparse-view inputs into long-range, high-fidelity novel-view sequences; Difix3D+ [73] applies a single-step diffusion enhancer to rendered novel views to correct artifacts in underconstrained regions and distill the improvements back into NeRF/3DGS representations; and TrajectoryCrafter [87] formulates camera-controllable video generation for monocular videos as trajectory redirection, conditioning a dual-stream diffusion backbone on point-cloud renderings and source frames to follow user-specified camera paths. Despite their strong spatial–temporal consistency and viewpoint controllability, these reconstruction-based approaches are mostly tailored to static or quasi-static scenes and rely on curated data or heavyweight offline reconstruction, limiting scalability to in-the-wild monocular videos.
@@ 3.1 Pose-Free Feed-Forward 4DGS Reconstruction
Our feed-forward model is partially built upon VGGT [67] backbone. For simplicity, we mainly introduce how we make VGGT dynamic and “Gaussianized”.
@@
Given a monocular video $\{\boldsymbol{I}_{t}\in\mathbb{R}^{H\times W\times 3}\}_{t=1}^{T}$, VGGT extracts the frame-wise features using the pretrained DINOv2 [54]. These features, concatenated with camera tokens and register tokens, are fed into a series of Alternating-Attention blocks [67], obtaining so-called frame features. While this process effectively aggregates spatial information, they are insufficient for motion modeling due to temporal unawareness.
We introduce a bidirectional motion-encoding branch. Different from uni-directional motion in 4DGT [78], the bidirectional prediction distinguishes the instantaneous velocity between $t\rightarrow t+1$ and $t\rightarrow t-1$. Such a distinction facilitates temporal Gaussian interpolation between two consecutive timestamps.
@@
where each Gaussian $i$ is parameterized by: 3D position $\boldsymbol{\mu}_{i}$, opacity $\boldsymbol{\alpha}_{i}$, rotation $\boldsymbol{r}_{i}$, scale $\boldsymbol{s}_{i}$, and spherical harmonics coefficients $\boldsymbol{sh}_{i}$, as inherited from 3D Gaussians [36]. For bidirectional motion modeling, we introduce forward and backward velocities $\boldsymbol{v}^{+}_{i}$, $\boldsymbol{v}^{-}_{i}$, and forward and backward angular velocities $\boldsymbol{\omega}^{+}_{i}$, $\boldsymbol{\omega}^{-}_{i}$. In addition, we adopt a life span $\boldsymbol{\tau}_{i}$ following the common practice in 4DGS.
@@ 3.2 Reconstruction-Guided Video Generation
Our generation model is expected to generate high-quality novel views from low-quality novel view renderings, necessitating such training pairs. For multi-view or static datasets [85, 44], we can easily get such training pairs as in ViewCrafter [88]. However, for in-the-wild monocular videos, we need to carefully simulate degradation renderings paired with ground-truth monocular frames. Therefore, we propose three techniques to simulate the degradation rendering patterns based on monocular videos.
@@
(2) Average Geometry Filter for flying-edge-pixel and distortion simulation. In addition to the occlusion, another typical degradation pattern is the flying pixels in depth-discontinuous edges. The network has tendency to produce average depth value at those edges to minimize regression loss, as also confirmed by [76]. From a first-principles perspective, we propose to use a average filter to create such averaged depth patterns. Specifically, we render depth in the transformed novel trajectory and apply an average filter in the rendered depth map. We then adjust the center position of each Gaussian according to the average filtered depth value. When such modified Gaussians are rendered back into the original views, the flying-pixel pattern appears as shown in Fig. 3 (b). We can further apply a larger filter kernel to simulate spatially broader distortions shown in Fig. 3 (c), caused by potential depth error.
@@
We use the obtained degraded renderings as conditions for generation and the original videos as targets. The rendered conditions include multiple modalities, including RGB images, depth maps, and masks binarized from opacity maps to indicate the empty regions. Plüker embeddings of the original trajectory are also computed to provide explicit 3D camera motion information [9]. We introduce a control branch to incorporate them into the generation model like [90, 33, 30, 80]. During training, we only train the control branch while freezing the video generation model, not only for training efficiency, but more importantly, to make NeoVerse accessible to powerful distillation LoRAs [24] to speed up the generation process.
@@ 3.3 Training Scheme
where $\mathcal{L}_{\text{rgb}}$ is the photometric loss between rendered and ground-truth images, including an $L_{2}$ loss and LPIPS [91] loss. The camera loss $\mathcal{L}_{\text{camera}}$ and depth loss $\mathcal{L}_{\text{depth}}$ supervise the predicted camera parameters and depth maps following VGGT [67]. Notably, $\mathcal{L}_{\text{depth}}$ also contains the supervision for rendered depth from Gaussians. The motion loss $\mathcal{L}_{\text{motion}}=\sum_{i}\|\hat{\boldsymbol{v}}^{+}_{i}-\boldsymbol{v}^{+}_{i}\|+\|\hat{\boldsymbol{v}}^{-}_{i}-\boldsymbol{v}^{-}_{i}\|$ adds supervision on the predicted bidirectional velocities, where $\hat{\boldsymbol{v}}^{+}_{i}$ and $\hat{\boldsymbol{v}}^{-}_{i}$ are the ground-truth forward and backward velocities computed from some dynamic 3D datasets [93, 34, 21, 53, 63, 8]. To prevent the Gaussians from becoming erroneously transparent, we introduce a regularization loss $\mathcal{L}_{\text{regular}}=\sum_{i}|1-\boldsymbol{A}_{i}|$, where $\boldsymbol{A}_{i}$ is rendered accumulated opacity map.
@@
For generation model training, we adopt Rectified Flow [16] and Wan-T2V [65] 14B to model the denoising diffusion process. The whole training process is performed on monocular videos. Given a monocular video, we first utilize on-the-fly reconstruction from sparse key frames to obtain 4DGS and simulate degradation renderings as conditions $c_{\text{render}}$. For the video latent $x_{1}$ and sampled noise $x_{0}\sim\mathcal{N}(0,I)$, the training objective of generation model $f_{\theta}$ is formulated as
@@
where $x_{t}$ is a linear interpolation between $x_{1}$ and $x_{0}$ at timestamp $t$, $v_{t}=x_{1}-x_{0}$ is ground-truth velocity. $c_{\text{text}}$ is the text condition extracted from the video caption using a language model like umT5 [14]. Renderings $c_{\text{render}}$ are input into the generation model through a control branch like [90, 33].
@@ 4.1 Implementation
For reconstruction, we follow the learning rate schedule of VGGT [67]. We resize all input videos to have a longest edge of 560 pixels. GSplat [84] is adopted as the Gaussian Splatting rendering backend. For the generation, the video resolution is fixed at $336\times 560$ and the length is set to $81$ frames. The training is conducted on 32 A800 GPUs, where the first stage trains 150K iterations and the second stage trains 50K iterations. More training details can be found in the supplementary material.
@@
We collect 18 public datasets following CUT3R [68], including Arkitscenes [4], DL3DV [44], PointOdyssey [93], Kubric [21], Waymo [63], SpatialVID [66], GFIE [27], etc. Besides the above datasets, we further curate a large-scale self-collected monocular video dataset from the internet, containing over 1M videos from diverse scenarios. More details about datasets are provided in the supplementary material.
@@ 4.2 Quantitative Evaluation
Our reconstruction results on both static and dynamic datasets are shown in Table 1 and Table 2, respectively. Our reconstruction part achieves state-of-the-art performance among all metrics. Recent reprints MoVieS [42] and StreamSplat [74] are not listed in the table because they are neither open-sourced nor provide a detailed evaluation protocol. Our detailed evaluation protocols are provided in the supplementary material.
Method VRNeRF [77] (16 views) Scannet++ [85] (32 views)
@@
NoPoSplat [83] 11.27 0.408 0.620 8.69 0.312 0.614
Flare [92] 12.62 0.597 0.623 12.19 0.619 0.611
AnySplat [32] 18.02 0.705 0.366 22.79 0.773 0.217
@@
In Table 3, we compare the generation performance with related work TrajectoryCrafter [87] and ReCamMaster [2], demonstrating better performance. We conduct more analysis in the section of qualitative evaluation.
@@
Table 3 also shows the efficiency evaluation of both the reconstruction stage and the generation stage. Thanks to our intentional design of condition injection in Sec. 3.2, our generation process gets significantly accelerated by the off-the-shelf distillation technique [15]. More importantly, as discussed in Sec. 3.2, our bidirectional motion design enables more efficient reconstruction from sparse key frames without loss of generation performance.
@@
TrajectoryCrafter [87] 49 25 121 146 83.02 88.58 94.71 97.64 44.63 54.59
ReCamMaster [2] 81 - 168 168 88.21 91.60 96.56 98.86 44.29 58.87
@@
Table 3: VBench [31] results for novel view generation. We randomly collect 100 unseen in-the-wild videos, each with 4 different camera trajectories, resulting in a total of 400 test cases. For a fair comparison of inference time, we resize all videos to $336\times 560$ resolution and report the average results over all test cases. The runtime evaluation is conducted on an A800 GPU.